# SmartSustain Recommender System: Navigating Sustainability Trade-offs in Personalized City Trip Planning

### v1

@@ Abstract.
Tourism Recommender Systems, Sustainable Tourism, User-Centric Design, Web Application
copyright: Nonejournalyear: 2025conference: Workshop on Recommender Systems for Sustainable Development (RS4SD), co-located with CIKM’2025; November 10–14, 2025; Seoul, Republic of Korea
1. Introduction
Online travel platforms have democratized access to global destinations; yet, their recommendation algorithms have traditionally optimized for user preferences and cost, largely ignoring the substantial environmental and societal costs associated with tourism (Banerjee, 2023). As awareness grows around issues like climate change and over-tourism, there is an urgent need for tools that not only present sustainable options but also help users comprehend the complex trade-offs involved. A more sustainable travel choice may involve a higher price, a longer journey, or a destination that is a slightly weaker match for a user’s stated interests.
@@ 1. Introduction
Our work builds on existing research in sustainable TRS, which has sought to integrate environmental considerations into travel planning (Banerjee et al., 2024a, b). However, many of these systems focus primarily on providing information about sustainable options without adequately addressing the cognitive and emotional barriers that users face when making travel decisions (Banerjee et al., 2024c). In contrast, the SmartSustain Recommender emphasizes user engagement through interactive visualizations and real-time feedback, making the trade-offs between sustainability and personal preferences more transparent and manageable. Additionally, our system extends the Green Destination Recommender (Banerjee et al., 2024c) by providing a more personalized and context-aware approach to guiding users toward eco-friendly travel choices. A preliminary user study with 21 participants further confirmed the system’s usability and perceived effectiveness in promoting sustainability, while also highlighting opportunities to refine certain interface elements.
@@ 2. Related Work
Recent research has begun to integrate sustainability into tourism recommenders. The work of Banerjee et al. (2024a, c) forms the foundation for our system by proposing a framework to model sustainable city trips using CO2e emissions, popularity, and seasonality. Other research has focused on multi-stakeholder fairness, balancing the needs of travelers, local businesses, and residents to mitigate issues like over-tourism (Abdollahpouri and Burke, 2021; Merinov and Ricci, 2024). Mauro et al. (2024) explored nudging users towards sustainable points-of-interest. Our work extends these efforts by focusing specifically on the communication of trade-offs to the end-user, moving beyond algorithmic optimization to user-centric interaction design.
@@
Making complex choices can lead to cognitive overload and decision fatigue (Echterhoff et al., 2024). Schnabel et al. (2016) demonstrated that tools like shortlists can reduce this burden by helping users manage options. We apply this principle by designing components that offload the mental effort of comparing multiple factors. Furthermore, research by Okoso et al. (2024) has shown that the tone of explanations can significantly impact user engagement. We incorporate this by using motivational and positive framing in our nudges, aiming to encourage sustainable choices rather than merely presenting data. The SmartSustain Recommender combines these user experience (UX) principles to create an environment where users do not feel burdened by too many choices.
Figure 1. Card view with sustainability tags, transport summary, and interest alignment bars. The dynamic banner below, highlighted in red, aims to encourage users to make more sustainable choices.
Figure 2. Map view with interactive markers and polylines representing destinations and transport options. Upon clicking on one of the markers, users can view detailed information about the destination and available transportation (highlighted in red).
3. User Journey and Interface
@@
Figure 3. Booking component with transport and accommodation options showing financial and environmental impacts.
@@
Figure 4. Different banners promoting sustainable travel options. They are contextually displayed based on user interactions and preferences to nudge them towards greener choices.
@@
Color coding facilitates rapid interpretation: green icons denote low-emission options, while red icons indicate high-impact choices, similar to the traffic light color coding scheme (Magazine, 2020). Labels such as “Best Match,” “Green Runner-up,” or “Eco-Bronze Medallist” provide comparative cues. We also use real-time animations, such as CO2 progress bars or tree icons, to visualize emission thresholds and make trade-offs tangible, creating an engaging and informative user experience.
Figure 5. Alternative destination suggestions shown on the booking page when a high-emission option is selected. These alternatives are greener and align well with the user’s interests, encouraging more sustainable choices.
@@
4. System Design and Validation
@@
4.1. Core Recommendation Logic
Destinations are ranked using a linear, interpretable score where lower values indicate better outcomes, preserving a consistent “less is better” interpretation across sustainability criteria. Our overall scoring function (Sustainability Score) is inspired from Banerjee et al. (2024a), where the authors use a weighted sum of multiple factors to rank travel destinations and assign a sustainability indicator to the destination. The score combines transport efficiency, popularity (in terms of the number of user-generated content, i.e., ratings and reviews at the destination), seasonality (defined in terms of the estimated crowdedness of the destination during the intended travel month), and interest alignment, along with personalization that reflects user-prioritized sustainability attributes such as air quality, climate vulnerability, and walkability. When personalization is absent, transport and interest alignment are automatically emphasized. For the start weights, we use the same weights as in the original work.
@@ 4.2. Implementation
4.3. Preliminary User Study

### v2

# Workshop on Recommender Systems for Sustainable Development (RS4SD), co-located with CIKM’2025, November 10–14,2025, Seoul, Republic of Korea

@@ Abstract
keywords:
[…]
1 Introduction
Online travel platforms have democratized access to global destinations; yet, their recommendation algorithms have traditionally optimized for user preferences and cost, largely ignoring the substantial environmental and societal costs associated with tourism [1]. As awareness grows around issues like climate change and over-tourism, there is an urgent need for tools that not only present sustainable options but also help users comprehend the complex trade-offs involved. A more sustainable travel choice may involve a higher price, a longer journey, or a destination that is a slightly weaker match for a user’s stated interests.
@@ 1 Introduction
Our work builds on existing research in sustainable TRS, which has sought to integrate environmental considerations into travel planning [2, 3]. However, many of these systems focus primarily on providing information about sustainable options without adequately addressing the cognitive and emotional barriers that users face when making travel decisions [4]. In contrast, the SmartSustain Recommender emphasizes user engagement through interactive visualizations and real-time feedback, making the trade-offs between sustainability and personal preferences more transparent and manageable. Additionally, our system extends the Green Destination Recommender [4] by providing a more personalized and context-aware approach to guiding users toward eco-friendly travel choices. A preliminary user study with 21 participants further confirmed the system’s usability and perceived effectiveness in promoting sustainability, while also highlighting opportunities to refine certain interface elements.
@@ 2 Related Work
Recent research has begun to integrate sustainability into tourism recommenders. The work of Banerjee et al. [2, 4] forms the foundation for our system by proposing a framework to model sustainable city trips using CO2e emissions, popularity, and seasonality. Other research has focused on multi-stakeholder fairness, balancing the needs of travelers, local businesses, and residents to mitigate issues like over-tourism [5, 6]. Mauro et al. [7] explored nudging users towards sustainable points-of-interest. Our work extends these efforts by focusing specifically on the communication of trade-offs to the end-user, moving beyond algorithmic optimization to user-centric interaction design.
@@
Making complex choices can lead to cognitive overload and decision fatigue [8]. Schnabel et al. [9] demonstrated that tools like shortlists can reduce this burden by helping users manage options. We apply this principle by designing components that offload the mental effort of comparing multiple factors. Furthermore, research by Okoso et al. [10] has shown that the tone of explanations can significantly impact user engagement. We incorporate this by using motivational and positive framing in our nudges, aiming to encourage sustainable choices rather than merely presenting data. The SmartSustain Recommender combines these user experience (UX) principles to create an environment where users do not feel burdened by too many choices.
Figure 1: Card view with sustainability tags, transport summary, and interest alignment bars. The dynamic banner below, highlighted in red, aims to encourage users to make more sustainable choices.
Figure 2: Map view with interactive markers and polylines representing destinations and transport options. Upon clicking on one of the markers, users can view detailed information about the destination and available transportation (highlighted in red).
3 User Journey and Interface
@@
Figure 3: Booking component with transport and accommodation options showing financial and environmental impacts.
@@
Figure 4: Different banners promoting sustainable travel options. They are contextually displayed based on user interactions and preferences to nudge them towards greener choices.
@@
Color coding facilitates rapid interpretation: green icons denote low-emission options, while red icons indicate high-impact choices, similar to the traffic light color coding scheme [11]. Labels such as “Best Match,” “Green Runner-up,” or “Eco-Bronze Medallist” provide comparative cues. We also use real-time animations, such as CO2 progress bars or tree icons, to visualize emission thresholds and make trade-offs tangible, creating an engaging and informative user experience.
Figure 5: Alternative destination suggestions shown on the booking page when a high-emission option is selected. These alternatives are greener and align well with the user’s interests, encouraging more sustainable choices.
@@
4 System Design and Validation
@@
4.1 Core Recommendation Logic
Destinations are ranked using a linear, interpretable score where lower values indicate better outcomes, preserving a consistent “less is better” interpretation across sustainability criteria. Our overall scoring function (Sustainability Score) is inspired from Banerjee et al. [2], where the authors use a weighted sum of multiple factors to rank travel destinations and assign a sustainability indicator to the destination. The score combines transport efficiency, popularity (in terms of the number of user-generated content, i.e., ratings and reviews at the destination), seasonality (defined in terms of the estimated crowdedness of the destination during the intended travel month), and interest alignment, along with personalization that reflects user-prioritized sustainability attributes such as air quality, climate vulnerability, and walkability. When personalization is absent, transport and interest alignment are automatically emphasized. For the start weights, we use the same weights as in the original work.
@@ 4.2 Implementation
4.3 Preliminary User Study
Figure 6: User feedback on the SmartSustain Recommender.